OpenStudy (anonymous):

Which set of side lengths shows similar triangles? Answer Triangle ABC : 32, 20, 32; Triangle XYZ: 30, 16, 16 Triangle ABC : 110, 80, 60; Triangle XYZ: 6, 8, 5.5 Triangle ABC : 40, 20, 50; Triangle XYZ: 10, 12.5, 10 Triangle ABC : 30, 30, 60; Triangle XYZ: 15, 15, 30

OpenStudy (amistre64):

similar triangles will have the same ratio of sides

OpenStudy (amistre64):

for example: ABC = 3,4,5 XYZ = 6,8,10 then \(\dfrac{ABC}{XYZ}=\dfrac{3,4,5}{6,8,10}=\dfrac{1}{2},\dfrac{1}{2},\dfrac{1}{2}\)

OpenStudy (amistre64):

or if the numbers are simple enough, one will be a scalar of the other